Find all solutions of the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to two decimal places where appropriate.)
4 cos θ + 1 = 0

Answer by lwsshak3(11628)

Find all solutions of the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to two decimal places where appropriate.)
4 cos θ + 1 = 0
cos θ=-1/4
reference angle≈1.32 in quadrants II and III where cos<0
θ≈1.82+2πk, 4.46+2πk, k=any integer